OpenStudy (anonymous):

when 27x^2z/-3x^2z^6 is completely simplified the exponent of variable z is?

OpenStudy (mathstudent55):

Since all you're concerned is with the exponent of z, let's just look at the z variables. You have \( \dfrac{z}{z^6} \) Remember the rule: \(\dfrac{a^m}{a^n} = a^{m - n} \) Also, remember that z is the same as \(z^1\), not \(z^0\).

OpenStudy (anonymous):

so it would be z^5?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You subtract the exponents, but the exponent of 1 is in the numerator, so it's \(\dfrac{z}{z^6} = z^{1 - 6} = z^{-5} = \dfrac{1}{z^5}\) If you write the result without a fraction, the exponent is -5. If you write the result as a fraction, then z^5 is in the denominator and the exponent is 5.